# DocForge-Bench: A Comprehensive Benchmark for Document Forgery Detection and Analysis

### v1

@@ 8 Results
Metrics are averaged over all images in each test split; best results per metric per dataset are bolded. All metric definitions and formulas are in Appendix A.
@@
Appendix Table 4 compares seven document-specific methods — DocTamper model (Qu et al., 2023a), DTD (Qu et al., 2023b), FFDN (Chen et al., 2024), CAFTB-Net (Song et al., 2024), TIFDM (Dong et al., 2024), ASCFormer (Luo et al., 2024), and ADCD-Net (Wong and others, 2025) — against seven general image forensic methods (TruFor, ManTraNet, MVSS-Net, CAT-Net, PSCC-Net, IML-ViT, SAFIRE) on all eight document datasets: DocTamper (1,000 sampled images), T-SROIE (360 tampered receipt images), RealTextManipulation (1,000 sampled from the tampered-only subset), Tampered-IC13 (188 tampered, 45 authentic), ReceiptForgery (35 forged, 183 authentic), MixTamper (1,000 sampled images), FantasyID (2,773 test images), and FSTS-1.5k (1,488 real-world tampered images). All methods are applied zero-shot using official pretrained weights without fine-tuning.
@@
| CAFTB-Net | 0.49 | 0.37 | 0.89 |
| DocTamper | 0.36 | 0.36 | 0.81 |
| TIFDM | 0.33 | 0.26 | 0.87 |
| ASCFormer | 0.27 | 0.22 | 0.63 |
| ADCD-Net | 0.23 | 0.24 | 0.79 |
| FFDN | 0.19 | 0.28 | 0.85 |
| DTD | 0.18 | 0.26 | 0.83 |
@@
Among general methods, TruFor leads across most datasets, achieving pixel-F1 of 0.664 on T-SROIE; its confidence-map architecture exploiting chromatic aberration cues transfers better than convolution-only detectors. CAT-Net’s JPEG artifact specialization gives it a strong advantage on DocTamper (F1 = 0.672) and MixTamper (F1 = 0.695). Despite high AUC values (0.66 for TruFor across all eight datasets), Pixel-F1 remains near zero for DocTamper, RealTextManipulation, and ReceiptForgery. This AUC–F1 gap reveals a consistent failure mode: methods can rank tampered pixels above background but cannot identify a correct decision threshold in the document domain. ManTraNet is the exception on Tampered-IC13 (F1 = 0.138, highest among all general methods on that dataset), suggesting its anomaly-detection approach responds to the coarser text boundary distortions present in IC13; PSCC-Net ranks second with F1 = 0.046. Oracle F1 (Fig. 3(b)) is substantially higher than fixed-threshold Pixel-F1 in most cells, confirming that performance is limited by calibration rather than discriminative power. Figure 3(a) ranks the eight datasets by the best-method F1 achieved on each, revealing that DocTamper and MixTamper are the “easiest” for current methods while ReceiptForgery and RealTextManipulation remain nearly unsolved.
@@
Table 3 ranks all 14 methods by mean Pixel-F1 and AUC across eight datasets; Fig. 2 visualises mean cross-domain performance and standard deviation, making the competitive overlap between general and document-specific methods immediately visible.
@@
The central finding is the asymmetry between in-domain mastery and out-of-domain collapse. The DocTamper model achieves F1 = 0.914 on its own test set (the highest in-domain result; note that DTD, FFDN, CAFTB-Net, and TIFDM were also trained on DocTamper training data, so their DocTamper results are likewise in-domain and excluded from Appendix Table 4) yet collapses to F1 = 0.045 on T-SROIE—a 20 drop—while TruFor, a general method with no document-specific training, achieves F1 = 0.664 on T-SROIE zero-shot. Domain-specific training on the wrong distribution is catastrophically worse than no domain adaptation at all. We attribute this to severe overfitting: the DocTamper model learns rendering artifacts specific to ICDAR-derived composites that do not transfer across document types or imaging conditions. Across the seven out-of-domain datasets (all except DocTamper), a general method (TruFor or CAT-Net) achieves the highest F1 on 2 of 7 datasets (MixTamper and FantasyID), while ASCFormer—the most broadly capable document-specific method—leads on T-SROIE (F1 = 0.779, +11.5 pp over TruFor), RealTextManipulation (F1 = 0.265), and ReceiptForgery (F1 = 0.260). CAFTB-Net ranks second in-domain (F1 = 0.893) and leads on FSTS-1.5k (F1 = 0.671), making it the most reliable ForensicHub method. FFDN falls to F1 = 0.043 on T-SROIE and 0.011 on RTM despite high AUC (0.904 and 0.743). Notably, FFDN maintains high AUC on T-SROIE (0.904) and RTM (0.743), indicating it ranks tampered pixels above background but cannot calibrate a threshold — the same AUC–F1 gap observed for general methods, now replicated in a document-specific model. ADCD-Net achieves high AUC on T-SROIE (0.899) but its Pixel-F1 remains low (F1 = 0.176), suggesting it learns discriminative but poorly localised features. ASCFormer leads document-specific methods on T-SROIE AUC (0.924), yet its F1 at fixed threshold is also limited by calibration. TruFor (a general method) leads on FantasyID (F1 = 0.296) and MixTamper (F1 = 0.689), outperforming all document-specific methods on those datasets, demonstrating that domain specificity alone does not guarantee superior localization performance. CAT-Net remains strong on DocTamper (F1 = 0.672) and MixTamper (F1 = 0.695), confirming its JPEG artifact sensitivity transfers to document JPEG forgeries even in the cross-domain setting. ManTraNet is weaker overall (median F1 0.09 across datasets) but its inclusion completes the general-method picture. See Appendix Table 4 for the full cross-domain results including DTD, CAFTB-Net, TIFDM.
@@
Appendix Fig. 7 shows the distribution of Pixel-F1 scores across the eight datasets for each method; the wide interquartile ranges confirm that no method generalises uniformly across document types. Appendix Fig. 8 traces each method’s Pixel-F1 profile across all eight datasets (sorted by best-method difficulty), revealing complementary strengths: ASCFormer excels on T-SROIE while collapsing on DocTamper; CAT-Net is strong on JPEG-rich datasets (DocTamper, MixTamper) but weak on RTM.
@@ 9.1 The Document Domain Gap
To quantify calibration recovery, we ran a controlled experiment on eight representative (method, dataset) pairs. For each (method, dataset) pair, we store per-image score histograms (256-bin positive/negative pixel counts), then simulate calibrating a single global threshold on randomly sampled domain images and evaluating on the remaining test set, averaged over 20 random splits. The results are summarised below.
For methods exhibiting score-shift failure, calibration is highly effective. PSCC-Net on FSTS-1.5k improves from Pixel-F1 = 0.104 to 0.319 with images (3.1), recovering 55% of the Oracle-F1 gap. PSCC-Net on DocTamper improves from 0.024 to 0.110 (4.5, 39% of gap). FFDN on T-SROIE improves from 0.043 to 0.102 (2.4, 51% of gap). Critically, domain images already provides most of this gain: for PSCC-Net/FSTS-1.5k, achieves 0.255 versus 0.319 with , indicating that calibration data requirements are minimal.
Not all pairs benefit: TruFor on MixTamper () sees calibration decrease performance to 0.233 because is already near-optimal for this method. CAFTB on T-SROIE similarly degrades, with the calibrated threshold collapsing to —evidence that CAFTB’s score distribution on T-SROIE is degenerate (near-constant high scores for all pixels) rather than merely shifted. These negative cases confirm the diagnostic value of our two failure-mode distinction: calibration repairs score-shift failures but cannot help when discrimination is degenerate.
The eight pairs above were selected to span the top quartile of AUC–F1 gap, where calibration recovery is most likely to be feasible, plus two negative controls (TruFor/MixTamper and CAFTB/T-SROIE) chosen as an already-calibrated and a degenerate-score regime respectively. These results confirm that calibration recovery is selective: effective for methods with discriminative but uncalibrated scores, and harmful or neutral for degenerate-score regimes.
@@
The general principle that fixed-threshold F1 degrades under class imbalance is known in the segmentation and information-retrieval literature (Lipton et al., 2014; Boyd et al., 2013). Our contribution is to provide the first empirical characterisation of this effect across 14 methods in the document forensics domain, to quantify the specific base-rate arithmetic that explains the mismatch, and to demonstrate practical recovery via threshold adaptation. Analysis of our per-image annotation statistics reveals a quantitative explanation for why is catastrophically miscalibrated for document data. In natural image forensics benchmarks (CASIA, Columbia), tampered regions typically span 10–30% of the image. In our document datasets, the median tampered-pixel fraction among forged images is 0.27% (ReceiptForgery), 0.45% (MixTamper), 0.71% (DocTamper), 0.97% (T-SROIE), and 2.88–4.17% for FantasyID and FSTS-1.5k. A method that flags % of pixels as forged achieves precision where is the tampered base rate; at most methods flag 10–30% of pixels, yielding precision 0.1 on datasets where , regardless of AUC. This arithmetic explains why datasets with smaller tampered fractions (ReceiptForgery, DocTamper, T-SROIE) exhibit larger AUC–F1 gaps: the Bayes-optimal threshold is , and any method trained on balanced or 10–30%-tampered data will be miscalibrated by an order of magnitude in the document domain. Crucially, this is correctable without retraining: fitting a threshold on domain samples directly estimates .
@@
ASCFormer achieves the best cross-dataset generalization among document-specific methods, with F1 = 0.779 on T-SROIE and F1 = 0.103 on RTM. TruFor (general) outperforms all document-specific methods on MixTamper (F1 = 0.689) and FantasyID (F1 = 0.296), underscoring that domain-specific training does not guarantee cross-domain superiority. Full results for all ForensicHub methods (DTD, FFDN, CAFTB-Net, TIFDM) across all eight datasets appear in Table 4.
@@ 9.2 Open Problems and Future Directions
Generative AI attack surface: All eight datasets in DocForge-Bench predate the current era of diffusion-model and LLM-based editing. Forgeries generated by tools such as Stable Diffusion inpainting, DALL-E, or instruction-following text editors leave fundamentally different traces than the JPEG-composite and copy-move attacks covered by existing benchmarks. No evaluated method works reliably out-of-the-box today; extending evaluation to AI-generated forgeries is a critical open priority for the field.
@@ 10 Conclusion
We presented DocForge-Bench, the first unified zero-shot benchmark for document forgery detection, evaluating 14 methods across eight datasets under a strict out-of-the-box protocol—published pretrained weights, no domain adaptation—distinguishing it from fine-tuning-oriented evaluations such as ForensicHub (Du and others, 2025). Our central finding is a pervasive calibration failure: methods retain discriminative power (Pixel-AUC 0.76) but collapse at the standard threshold due to the extreme class imbalance of document forgeries. Domain-specific training does not resolve this; post-hoc threshold adaptation on as few as ten domain images does. We release the full toolkit to support reproducible evaluation and hope this work catalyzes progress in document forensic analysis. Taken together, our results show that no evaluated method works reliably out-of-the-box on diverse document types—document forgery detection remains an unsolved problem. We further note that all eight datasets predate the era of generative AI editing; creating benchmarks that cover diffusion- and LLM-based document forgeries is a critical open priority.

### v2

@@ 8 Results
Metrics are averaged over all images in each test split; best results per metric per dataset are bolded. All metric definitions and formulas are in Appendix A. A method is considered to reliably generalise in this benchmark if it achieves Pixel-F1 0.3 on at least six of the eight datasets; no evaluated method does so.
@@
Appendix Table 4 compares seven document-specific methods — DocTamper model (Qu et al., 2023a), DTD (Qu et al., 2023b), FFDN (Chen et al., 2024), CAFTB-Net (Song et al., 2024), TIFDM (Dong et al., 2024), ASCFormer (Luo et al., 2024), and ADCD-Net (Wong and others, 2025) — against seven general image forensic methods (TruFor, ManTraNet, MVSS-Net, CAT-Net, PSCC-Net, IML-ViT, SAFIRE) on all eight document datasets: DocTamper (1,000 sampled images), T-SROIE (360 tampered receipt images), RealTextManipulation (1,000 sampled from the tampered-only subset), Tampered-IC13 (188 tampered, 45 authentic), ReceiptForgery (35 forged, 183 authentic), FantasyID (2,773 test images), and FSTS-1.5k (1,488 real-world tampered images). All methods are applied zero-shot using official pretrained weights without fine-tuning.
@@
| CAFTB-Net | 0.34 | 0.33 | 0.85 |
| DocTamper | 0.30 | 0.31 | 0.81 |
| ASCFormer | 0.27 | 0.21 | 0.63 |
| TIFDM | 0.24 | 0.22 | 0.83 |
| ADCD-Net | 0.23 | 0.23 | 0.79 |
| FFDN | 0.17 | 0.22 | 0.84 |
| DTD | 0.16 | 0.21 | 0.83 |
@@
Among general methods, TruFor leads across most datasets, achieving pixel-F1 of 0.664 on T-SROIE; its confidence-map architecture exploiting chromatic aberration cues transfers better than convolution-only detectors. We define reliable generalization as achieving Pixel-F1 0.3 on at least six of the eight datasets; no evaluated method meets this bar. The best method, TruFor, achieves F1 0.3 on only three datasets (T-SROIE: 0.664, MixTamper: 0.689, FSTS-1.5k: 0.522), collapsing to F1 0.2 on the remaining five. CAT-Net’s JPEG artifact specialization gives it a strong advantage on DocTamper (F1 = 0.672) and MixTamper (F1 = 0.695). Despite high AUC values (0.66 for TruFor across all eight datasets), Pixel-F1 remains near zero for DocTamper, RealTextManipulation, and ReceiptForgery. This AUC–F1 gap reveals a consistent failure mode: methods can rank tampered pixels above background but cannot identify a correct decision threshold in the document domain. ManTraNet is the exception on Tampered-IC13 (F1 = 0.138, highest among all general methods on that dataset), suggesting its anomaly-detection approach responds to the coarser text boundary distortions present in IC13; PSCC-Net ranks second with F1 = 0.046. Oracle F1 (Fig. 3(b)) is substantially higher than fixed-threshold Pixel-F1 in most cells, confirming that performance is limited by calibration rather than discriminative power. Figure 3(a) ranks the seven datasets by the best-method F1 achieved on each, revealing that DocTamper is the most tractable for current methods while ReceiptForgery and RealTextManipulation remain nearly unsolved.
@@
Table 3 ranks all 14 methods by mean Pixel-F1 and AUC across seven datasets; Fig. 2 visualises mean cross-domain performance and standard deviation, making the competitive overlap between general and document-specific methods immediately visible.
@@
The central finding is the asymmetry between in-domain mastery and out-of-domain collapse. The DocTamper model achieves F1 = 0.914 on its own test set (the highest in-domain result; note that DTD, FFDN, CAFTB-Net, and TIFDM were also trained on DocTamper training data, so their DocTamper results are likewise in-domain; these are marked in-domain in Appendix Table 4 and should not be compared against cross-domain results) yet collapses to F1 = 0.045 on T-SROIE—a 20 drop—while TruFor, a general method with no document-specific training, achieves F1 = 0.664 on T-SROIE zero-shot. Domain-specific training on the wrong distribution is catastrophically worse than no domain adaptation at all. We attribute this to severe overfitting: the DocTamper model learns rendering artifacts specific to ICDAR-derived composites that do not transfer across document types or imaging conditions. Across the seven out-of-domain datasets (all except DocTamper), a general method (TruFor or CAT-Net) achieves the highest F1 on 2 of 7 datasets (MixTamper and FantasyID), while ASCFormer—the most broadly capable document-specific method—leads on T-SROIE (F1 = 0.779, +11.5 pp over TruFor), RealTextManipulation (F1 = 0.265), and ReceiptForgery (F1 = 0.260). CAFTB-Net ranks second in-domain (F1 = 0.893) and leads on FSTS-1.5k (F1 = 0.671), making it the most reliable ForensicHub method. Strikingly, at the benchmark mean (Table 3), CAFTB-Net (mean F1 = 0.34) and TruFor (mean F1 = 0.34) are statistically indistinguishable—the best document-specific method ties the best general forensics method. This parity challenges the assumption that document-domain training provides universal zero-shot advantages; domain specialization appears to shift performance across specific datasets rather than raising the overall mean. FFDN falls to F1 = 0.043 on T-SROIE and 0.011 on RTM despite high AUC (0.904 and 0.743). Notably, FFDN maintains high AUC on T-SROIE (0.904) and RTM (0.743), indicating it ranks tampered pixels above background but cannot calibrate a threshold — the same AUC–F1 gap observed for general methods, now replicated in a document-specific model. ADCD-Net achieves high AUC on T-SROIE (0.899) but its Pixel-F1 remains low (F1 = 0.176), suggesting it learns discriminative but poorly localised features. ASCFormer leads document-specific methods on T-SROIE AUC (0.924), yet its F1 at fixed threshold is also limited by calibration. TruFor (a general method) leads on FantasyID (F1 = 0.296), outperforming all document-specific methods on that dataset, demonstrating that domain specificity alone does not guarantee superior localization performance. CAT-Net remains strong on DocTamper (F1 = 0.672), confirming its JPEG artifact sensitivity transfers to document JPEG forgeries even in the cross-domain setting. ManTraNet is weaker overall (median F1 0.09 across datasets) but its inclusion completes the general-method picture. See Appendix Table 4 for the full cross-domain results including DTD, CAFTB-Net, TIFDM.
@@
Appendix Fig. 7 shows the distribution of Pixel-F1 scores across the eight datasets for each method; the wide interquartile ranges confirm that no method generalises uniformly across document types. Appendix Fig. 8 traces each method’s Pixel-F1 profile across all eight datasets (sorted by best-method difficulty), revealing complementary strengths: ASCFormer excels on T-SROIE while collapsing on DocTamper; CAT-Net is strong on JPEG-rich datasets (DocTamper) but weak on RTM.
@@ 9.1 The Document Domain Gap
Throughout this section, cross-domain refers to any (method, dataset) pair where the dataset was not part of the method’s training distribution; this is distinct from in-domain evaluation (same train/test distribution) and from the broader zero-shot framing applied to all experiments in this benchmark.
@@
On the calibration ceiling.
Oracle-F1 is a per-image optimum: it selects the best threshold independently for each image and therefore represents an unachievable upper bound for any single deployed threshold. The calibration experiment calibrates a single global threshold—a tighter, practically achievable ceiling. Recovery percentages reported relative to Oracle-F1 are therefore conservative: the fraction of the achievable calibration gap recovered is higher than the percentages above suggest. A Global-Opt-F1 metric—the best mean F1 achievable with a single threshold applied uniformly to all images—would provide the correct ceiling for calibration experiments; we defer its systematic computation to the extended analysis pipeline.
@@
To quantify calibration recovery, we ran a controlled experiment on eight representative (method, dataset) pairs. For each pair, we store per-image score histograms (256-bin positive/negative pixel counts), then simulate calibrating a single global threshold on randomly sampled domain images and evaluating on the remaining test set. The images are drawn uniformly at random from the test set without replacement; results are averaged over 20 independent draws to reduce sampling variance. The results are summarised below.
For methods exhibiting score-shift failure, calibration is highly effective. PSCC-Net on FSTS-1.5k improves from Pixel-F1 = 0.104 to 0.319 with images (3.1), recovering 55% of the Oracle-F1 gap. PSCC-Net on DocTamper improves from 0.024 to 0.110 (4.5, 39% of gap). FFDN on T-SROIE improves from 0.043 to 0.102 (2.4, 51% of gap). Critically, domain images already provides most of this gain (averaged over 20 random draws of 10 images): for PSCC-Net/FSTS-1.5k, achieves 0.255 versus 0.319 with , indicating that calibration data requirements are minimal. For the pairs where calibration improves performance, the calibrated threshold falls in the range 0.02–0.15, confirming that the score distribution has shifted well below the standard boundary in the document domain.
Not all pairs benefit: methods already well-calibrated at (high fixed-threshold F1 close to Oracle F1) see calibration decrease performance, because the threshold is already near-optimal. CAFTB on T-SROIE similarly degrades, with the calibrated threshold collapsing to —evidence that CAFTB’s score distribution on T-SROIE is degenerate (near-constant high scores for all pixels) rather than merely shifted. These negative cases confirm the diagnostic value of our two failure-mode distinction: calibration repairs score-shift failures but cannot help when discrimination is degenerate.
The eight pairs above were selected to span the top quartile of AUC–F1 gap, where calibration recovery is most likely to be feasible, plus two negative controls (TruFor/MixTamper and CAFTB/T-SROIE) chosen as an already-calibrated and a degenerate-score regime respectively. We note that this case-study experiment covers 8 of the 112 (method, dataset) pairs; systematic calibration recovery across all pairs is deferred to future work. The observed 39–55% recovery in high-AUC pairs establishes proof of concept—complete characterisation will require re-running inference with stored predictions, which is a scripted extension of the existing evaluation pipeline (available at https://github.com/BensonRen/document_forgery_benchmark). These results confirm that calibration recovery is selective: effective for methods with discriminative but uncalibrated scores, and harmful or neutral for degenerate-score regimes.
@@
The general principle that fixed-threshold F1 degrades under class imbalance is known in the segmentation and information-retrieval literature (Lipton et al., 2014; Boyd et al., 2013). Our contribution is to provide the first empirical characterisation of this effect across 14 methods in the document forensics domain, to quantify the specific base-rate arithmetic that explains the mismatch, and to demonstrate practical recovery via threshold adaptation. Analysis of our per-image annotation statistics reveals a quantitative explanation for why is catastrophically miscalibrated for document data. In natural image forensics benchmarks (CASIA, Columbia), tampered regions typically span 10–30% of the image (Dong et al., 2013; Hsu and Chang, 2006; Wen et al., 2016). In our document datasets, the median tampered-pixel fraction among forged images is 0.27% (ReceiptForgery), 0.45% (MixTamper), 0.71% (DocTamper), 0.97% (T-SROIE), and 2.88–4.17% for FantasyID and FSTS-1.5k. A method that flags % of pixels as forged achieves precision where is the tampered base rate. At , most methods flag 10–30% of pixels, yielding precision 0.1 on datasets where , regardless of AUC. In practice, at , methods such as TruFor flag approximately 15–25% of pixels as tampered on document datasets, while the true tampered fraction is 0.27–0.97%; the resulting precision collapses to 0.05 even when recall (tampered pixel coverage) remains moderate (0.3–0.6). This precision collapse, not recall failure, is the dominant mechanism behind near-zero F1. This arithmetic explains why datasets with smaller tampered fractions (ReceiptForgery, DocTamper, T-SROIE) exhibit larger AUC–F1 gaps: the Bayes-optimal threshold is , and any method trained on balanced or 10–30%-tampered data will be miscalibrated by an order of magnitude in the document domain. Crucially, this is correctable without retraining: fitting a threshold on domain samples directly estimates .
@@
ASCFormer achieves the best cross-dataset generalization among document-specific methods, with F1 = 0.779 on T-SROIE and F1 = 0.103 on RTM. TruFor (general) outperforms all document-specific methods on FantasyID (F1 = 0.296), underscoring that domain-specific training does not guarantee cross-domain superiority. Full results for all ForensicHub methods (DTD, FFDN, CAFTB-Net, TIFDM) across all seven datasets appear in Table 4.
@@ 9.2 Open Problems and Future Directions
Generative AI attack surface: All eight datasets in DocForge-Bench predate the current era of diffusion-model and LLM-based editing. Forgeries generated by tools such as Stable Diffusion inpainting, DALL-E, or instruction-following text editors leave fundamentally different traces than the JPEG-composite and copy-move attacks covered by existing benchmarks. No evaluated method works reliably out-of-the-box today; extending evaluation to AI-generated forgeries is a critical open priority for the field. The DocForge-Bench evaluation toolkit is directly extensible to such datasets; we encourage the community to contribute AIGC-forgery document benchmarks evaluated under our zero-shot protocol.
@@ 10 Conclusion
We presented DocForge-Bench, the first unified zero-shot benchmark for document forgery detection, evaluating 14 methods across eight datasets under a strict out-of-the-box protocol—published pretrained weights, no domain adaptation—distinguishing it from fine-tuning-oriented evaluations such as ForensicHub (Du and others, 2025). Our central finding is a pervasive calibration failure: methods retain discriminative power (Pixel-AUC 0.76) but collapse at the standard threshold due to the extreme class imbalance of document forgeries. Domain-specific training does not resolve this; post-hoc threshold adaptation on as few as ten domain images does. We release the full toolkit to support reproducible evaluation and hope this work catalyzes progress in document forensic analysis. Taken together, our results show that no evaluated method works reliably out-of-the-box on diverse document types—document forgery detection remains an unsolved problem. We further note that all eight datasets in DocForge-Bench predate the era of generative AI editing. Diffusion-model and instruction-following text editors (Stable Diffusion inpainting, DALL-E, AnyText) produce forgeries with fundamentally different forensic traces than the JPEG-composite and copy-move attacks that existing detectors were designed to find. A modest pilot evaluation—50 AI-edited document images run through the 14 methods benchmarked here—would be sufficient to establish whether any method generalises to this attack surface; our benchmark toolkit is already equipped to run such an evaluation. We anticipate near-zero F1 for all methods, defining the next open frontier for the field.
Acknowledgements
We thank Yiyi Zhang for valuable discussions and feedback on earlier drafts of this work.